# DynamicControl : Adaptive Condition Selection for Improved Text-to-Image Generation

### v2

@@ 3.4 Multi-Control Adapter
After acquiring the well-pretrained condition evaluator, its robust understanding capabilities can be leveraged to score all input conditions. From the pool of scored conditions, only those that meet or exceed a predefined threshold are selected to participate in the subsequent optimization of the T2I model. This selective approach ensures that only the most relevant and high-quality conditions contribute to the training process, potentially enhancing the effectiveness and efficiency of the T2I model. Regarding the threshold setting, it is not manually predefined nor maintained consistently across all data pairs within the training set. Instead, it is configured as a parameter that is subject to learning, allowing the model to adaptively determine and adjust the threshold for various datasets. Consequently, this adaptive mechanism results in dynamic and diverse control conditions with no conflicts, both in quantity and type. These conditions are employed in the training process depending on the specific characteristics of each dataset. This approach ensures that the training is tailored to the unique demands and nuances of various data inputs.
As illustrated in Fig. 3, these selected conditions are then consumed by Mixture-of-Experts (MOEs) [50], where all conditions are captured in parallel as features of various low-level visual conditions. Subsequently, the extracted features are channeled into $L$ blocks, each comprising criss-cross attention [20] and cross attention mechanisms. Within each block, the input features are initially processed through $n(n\leq N)$ criss-cross attentions operating in parallel, aligning them across various feature space dimensions. Following this alignment, the diversified features are concatenated, which are then forwarded to the cross attention module. This sequential processing ensures a comprehensive integration of feature dimensions, enhancing the depth and relevance of the attention mechanisms applied. Finally, these multi-control embeddings are utilized to direct the generative process, ensuring that the output is aligned with the nuanced requirements specified by the multiple control conditions. For example, Fig. 4 shows the results of the six non-conflicting control conditions selected by our adapter dynamically. The conditions with high scores are added from left to right. It can be seen that as the high-scoring control conditions are gradually added, the generated results are of higher quality and gradually closer to the text description.
@@ 3.5 Training Strategy
The whole training of our network consists of two processes. The first process involves training the condition evaluator. The overall loss function, which combines LLM and diffusion losses, is represented as follows:
$\mathcal{L}_{\text{condi}}=\lambda_{1}\mathcal{L}_{\text{condition}}+\lambda_{2}\mathcal{L}_{\text{image}}+\lambda_{3}\mathcal{L}_{\text{LLM}}+\lambda_{4}\mathcal{L}_{\text{eval}},$ (10)
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ are positive constants to balance the different losses. This evaluator is then frozen, remaining unchanged during any subsequent optimization processes.
The second training process involves the multi-control diffusion model. As in Eq. 3, we define the training loss following LDM [44]. In our DynamicControl framework, akin to the scheme employed in ControlNet [63], we retain the pre-trained stable diffusion model in a fixed state while training its copied blocks.
@@ 4 Experiments
We validate the effectiveness of DynamicControl on five conditions with more common control conditions: canny, hed, segmentation mask, openpose and depth. Our evaluation primarily focuses on several leading methods in the realm of controllable text-to-image diffusion models, including Gligen [29], T2I-Adapter [35], ControlNet v1.1 [63], GLIGEN [29], Uni-ControlNet [65], UniControl [39], Cocktail [16] and ControlNet++ [27]. These methods are pioneering in their field and provide public access to their codes and model weights, which accommodate various image conditions. Although the models of other approaches such as AnyControl [54] are public, their code cannot be successfully run after many attempts. Implementation details including network structure, datasets, evaluation metrics, computational complexities, hyper-parameters of training and inference can be found in the supplementary material.
@@
Figure 5: Visualization comparison between official or re-implemented methods and our proposed model in different conditional controls with the same text prompt.
@@
Dataset T2I Model MultiGen-20M MultiGen-20M MultiGen-20M MultiGen-20M ADE20K COCO-Stuff
ControlNet [63] SDXL - - - 40.01 - -
T2I-Adapter [35] SDXL 28.03 - 63.89 39.76 - -
T2I-Adapter [35] SD1.5 23.66 - 60.17 48.40 12.60 -
Gligen [29] SD1.4 26.92 0.5641 69.88 38.82 23.77 -
Uni-ControlNet [65] SD1.5 27.31 0.6912 72.71 40.66 19.39 -
UniControl [39] SD1.5 30.83 0.7967 75.87 39.17 25.45 -
ControlNet [63] SD1.5 34.66 0.7622 - - 32.56 27.47
Cocktail [16] SD1.5 35.22 0.8152 78.82 35.90 36.55 29.68
ControlNet++ [27] SD1.5 37.04 0.8097 - 28.32 43.64 34.56
@@
Table 1: Controllability comparison under different conditional controls and datasets. We generate four groups of images and report the average result to reduce random errors.
Seg. Mask Canny Hed Openpose Depth
Gligen [29] 33.02 / 31.12 - 18.89 / 31.77 - 28.65 / 31.26 18.36 / 31.75
T2I-Adapter [35] 39.15 / 30.65 - 15.96 / 31.71 - 26.07 / 33.54 22.52 / 31.46
UniControlNet [65] 39.70 / 30.59 - 17.14 / 31.84 17.08 / 31.94 27.66 / 34.58 20.27 / 31.66
UniControl [39] 46.34 / 30.92 - 19.94 / 31.97 15.99 / 32.02 24.58 / 35.01 18.66 / 32.45
ControlNet [63] 33.28 / 31.53 21.33 / 13.31 14.73 / 32.15 15.41 / 32.33 - 17.76 / 32.45
Cocktail [16] 31.56 / 31.77 19.35 / 13.68 12.92 / 33.16 14.71 / 33.07 22.59 / 35.78 -
ControlNet++ [27] 29.49 / 31.96 19.29 / 13.13 18.23 / 31.87 15.01 / 32.05 - 16.66 / 32.09
Table 2: FID ($\downarrow$) / CLIP-score ($\uparrow$) comparison under different conditional controls and datasets. All the results are conducted on 512$\times$512 image resolution for fair comparisons. We generate four groups of images and report the average result to reduce random errors.
4.1 Quantitative Evaluation
[…]
Comparison of Image Quality. To ascertain whether enhanced controllability correlates with a reduction in image quality, we present the Fréchet Inception Distance (FID) metrics across multiple conditional generation tasks, as detailed in Tab. 2. We can find that our model quantitatively reveals superior performance across all conditions compared to existing approaches. This significant achievement indicates that DynamicControl effectively handles intricate combinations of multiple spatial conditions, producing high-quality, coherent outcomes that align well with the spatial conditions.
Comparison of CLIP Score. Our DynamicControl aims to boost the control over diffusion models by utilizing image-based conditions. To address concerns about its impact on text controllability, we employ CLIP-Score metrics to evaluate different methods across various datasets, ensuring that the generated images closely match the input text. As shown in Tab. 2, DynamicControl achieves superior CLIP-Score results on several datasets relative to existing methods. This indicates that our approach not only significantly improves conditional controllability, but also maintains the original model’s capability to generate images from text.
Figure 6: Comparison on Normal and Hedsketch conditions.
4.2 Qualitative Comparison
In Fig. 5 and Fig. 6, we visually compare different tasks — Canny, HED, Depth, Segmentation, Openpose, Normal, Hedsketch conditions. Our method consistently outperforms other models in terms of both visual quality and alignment with the specified conditions or prompts. This demonstrates the effectiveness of our approach in handling a diverse range of image generation tasks, while maintaining high fidelity to the input conditions. For instance, for the Hed and Segmentation task, the results generated by our model exhibit a higher degree of detail preservation and visual consistency. The outputs of our method maintain a faithful reproduction of the edge or mask information compared to other methods. In the Depth task, our model exhibits its superiority by producing plausible images characterized by seamless transitions and textures that mimic natural appearances. Furthermore, our model exhibits a more nuanced comprehension of geometric guidance pertaining to openpose and surface normals than the others. The outputs conditioned on openpose are noticeably more precise. The comparative visual analysis highlights the robustness and adaptability of our method, demonstrating its efficacy across a broad spectrum of tasks.
ADE20K MultiGen-20M
Base 28.18 38.76 31.26 22.58
+$\mathcal{L}_{\text{condition}}$ 30.11 33.59 33.05 19.62
+$\mathcal{L}_{\text{image}}$ 31.23 29.66 34.22 18.22
+$\mathcal{L}_{\text{LLM}}$ 33.04 27.58 35.26 16.85
+$\mathcal{L}_{\text{eval}}$ 34.38 25.23 36.68 15.28
Table 3: Ablation on loss functions. CLIP score and FID are reported on ADE20K and MultiGen-20M datasets.
4.3 Ablation Study
Loss Functions. To assess the effectiveness of different loss functions, we start with the base model which only contains the diffusion training loss. As shown in Tab. 3, adding the loss from the double-cycle controller significantly improves the alignment between the generated images and instruction conditions, as evidenced by the improved scores. Finally, the incorporation of combining the losses from LLM yields the best results across all metrics, underscoring the significance of condition evaluator for achieving high-fidelity and semantically accurate multiple condition generation.
Number of Conditions. As mentioned in Sec. 3.4, one of our key designs is to select dynamic number of conditions when performing the multi-control adapter, where we have tried different strategies of fixed numbers and adaptive number iteration. As shown in Fig. 7, in the scheme with a fixed number of iterations, the results vary with the iteration, but are still lower than those in the adaptive iteration method. This largely illustrates the effectiveness of the dynamic condition selection method in the multiple condition generation task.
5 Conclusion
In this paper, we demonstrate from both quantitative and qualitative perspectives that existing works focusing on controllable generation still fail to fully harness the potential of multiple control conditions, leading to inconsistency between generated images and input conditions. To address this issue, we introduce DynamicControl , it explicitly optimizes the consistency between multiple input conditions and generated images using an efficient condition evaluator to rank the conditions, which integrates MLLM’s reasoning capabilities into the T2I generation task. Furthermore, we also propose a novel and efficient multi-control adapter that selects different conditions adaptively, thus enabling dynamic multi-control alignment. Experimental results from various conditional controls reveal that DynamicControl substantially enhances controllability, without sacrificing image quality or image-text alignment. This provides fresh perspectives on controllable visual generation.
Supplementary Material
The supplementary material presents the following sections to strengthen the main manuscript:
  • Implementation details.
  • More experiments.
  • More qualitative results.
A1 Implementation Details
Datasets. In our framework, we include up to 12 control conditions, the majority of which are sourced from the MultiGen-20M dataset, as proposed by UniControl [39]. This dataset is a specialized subset derived from the larger LAION-Aesthetics [49]. More specifically, for the segmentation mask condition, our framework utilizes the ADE20K [66, 67] and COCOStuff [4] datasets, following ControlNet [63]. For instances where the text caption data is missing in ADE20K, we supplement it with data from ControlNet++ [27].
Evaluation Metrics. Our DynamicControl is trained using the training subsets of the respective datasets, and evaluations of all methods are conducted on the validation subsets. To ensure a fair comparison, the resolution for both training and inference in our framework is set at 512$\times$512 for all datasets and methods involved. Since the existing methods do not have so many control conditions, in order to facilitate fair comparison, we compare the quantitative evaluation on five conditions with more common control conditions: canny, hed, segmentation mask, openpose and depth. For each condition, controllability is assessed by quantifying the resemblance between the input conditions and the conditions extracted from the images generated by diffusion models. For the evaluation of segmentation, openpose and depth controls, we employ the mIoU, mAP and RMSE as metrics respectively, which is a common practice in related research fields. In the task of canny detection, we utilize the F1-Score as it effectively addresses the binary classification of pixels into categories of 0 (non-edge) and 1 (edge). This metric is particularly suitable given the pronounced long-tail distribution observed in edge data [59]. For evaluating other methods, we use the open-source code provided by the respective developers to generate images. We ensure a fair comparison by conducting evaluations under identical conditions, using the same datasets and without altering their inference configurations.
Baselines. Our evaluation primarily focuses on several leading methods in the realm of controllable text-to-image diffusion models, including Gligen [29], T2I-Adapter [35], ControlNet v1.1 [63], GLIGEN [29], Uni-ControlNet [65], UniControl [39], Cocktail [16] and ControlNet++ [27]. These methods are pioneering in their field and provide public access to their codes and model weights, which accommodate various image conditions. Other approaches such as AnyControl [54], although their models are public, their code cannot be successfully run after many attempts.
Training Details. During the first stage of training, we adopt the pre-trained LLaVAv1.1-7B [32] and QFormer [26] and employ DeepSpeed [1] Zero-2 to perform LoRA [15] fine-tuning. The Stable Diffusion-1.5 [44] is diffusion model pre-trained weights. The learning rate and weight decay parameters are set to 2e-4 and 0, respectively. In the second stage, the values of learning rate, weight decay, and warm-up ratio are set to 1e-5, 0, and 0.001, respectively. We train the model for totally 50K iterations. Furthermore, we take the AdamW [33] as the optimizer based on PyTorch Lightning [37] for both stages. Our full-version DynamicControl is trained on 8 H20 GPUs with the batch size of 4 and the whole training process can be completed within 3 days. It takes an average of 34s to generate one image on one H20 GPU. The values of $\lambda$ in Eq.10 are experimentally tested through a series of experiments, which are set to 1, 1, 2 and 1.5 respectively in our practice. These values vary from 0.1 to 3 at every 0.1 interval. In all the experiments, we adopt DDIM [51] sampler with 50 timesteps for all the compared methods.
A2 More Experiments
A2.1 Comparison on computational complexities

### v3

@@ 3.3 Multi-Control Adapter
After acquiring the well-pretrained condition evaluator, its robust understanding capabilities can be leveraged to score all input conditions. From the pool of scored conditions, only those that meet or exceed a predefined threshold are selected to participate in the subsequent optimization of the T2I model. This selective approach ensures that only the most relevant and high-quality conditions contribute to the training process, potentially enhancing the effectiveness and efficiency of the T2I model. Regarding the threshold setting, it is not manually predefined nor maintained consistently across all data pairs within the training set. Instead, it is configured as a parameter that is subject to learning, allowing the model to adaptively determine and adjust the threshold for various datasets. Consequently, as experimented in Fig. 7, this adaptive mechanism results in dynamic and diverse control conditions with no conflicts, both in quantity and type. These conditions are employed in the training process depending on the specific characteristics of each dataset. This approach ensures that the training is tailored to the unique demands and nuances of various data inputs.
As illustrated in Fig. 2, these selected conditions are then consumed by the following controllable image generation module. More details about can be found in the Appendix. Fig. 3 shows the results of the six non-conflicting control conditions selected by our adapter dynamically. The conditions with high scores are added from left to right. It can be seen that as the high-scoring control conditions are gradually added, the generated results are of higher quality and gradually closer to the text description.
@@ 3.4 Training Strategy
The whole training of our network consists of two processes. The first process involves training the condition evaluator, is represented as follows:
$\mathcal{L}_{\text{condi}}=\mathcal{L}_{\text{condition}}+\mathcal{L}_{\text{image}}+\lambda_{1}\mathcal{L}_{\text{LLM}}+\lambda_{2}\mathcal{L}_{\text{eval}},$ (5)
where $\lambda_{1}$ and $\lambda_{2}$ are positive constants to balance the different losses. This evaluator is then frozen, remaining unchanged during any subsequent optimization processes. The second training process involves the multi-control diffusion model.
@@ 4 Experiments
We validate the effectiveness of DynamicControl on five conditions with more common control conditions: canny, hed, segmentation mask, openpose and depth. Our evaluation primarily focuses on several leading methods in the realm of controllable text-to-image diffusion models, including Gligen (Li et al., 2023c), T2I-Adapter (Mou et al., 2024), ControlNet v1.1 (Zhang et al., 2023), GLIGEN (Li et al., 2023c), Uni-ControlNet (Zhao et al., 2024), UniControl (Qin et al., 2023), Cocktail (Hu et al., 2023) and ControlNet++ (Li et al., 2025). These methods are pioneering in their field and provide public access to their codes and model weights, which accommodate various image conditions. Although the models of other approaches such as AnyControl (Sun et al., 2024) are public, their code cannot be successfully run after many attempts. Implementation details including network structure, datasets, evaluation metrics, computational complexities, hyper-parameters of training and inference can be found in the Appendix C.
@@
Figure 4: Visualization comparison between selected SOTA methods and our proposed model in different conditional controls. Conditions are the conditions that are ultimately used by the model.
4.1 Main Results
Comparison in Multiple Conditions. Our DynamicControl aims to boost the control over diffusion models by multiple utilizing image-based conditions selection. As shown in Tab. 1, DynamicControl effectively addresses issues of line coarsening and subject distortion in multi - visual control conditions, as shown by quantitative (improved FID and leading MUSIQ scores) results. It also shows superior performance under full conditional control in terms of CLIP score, balancing visual control and text adherence better than other methods. Moreover, it maintains high consistency across different condition combinations and outperforms others in relevant metrics, especially under full-condition control.
Table 1: Comparison of generation quality, controllability, and text-image consistency on the MultiGen-20M (Li et al., 2025) and Subject-200K (Tan et al., 2024) datasets in multiple conditions. “all” under the Conditions column represents that all conditions are used for generation.
MultiGen-20M Subject-200K
Methods Conditions FID ($\downarrow$) CLIP Score ($\uparrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$) FID ($\downarrow$) CLIP Score ($\uparrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$)
T2I-Adapter (Mou et al., 2024) all 66.95 71.47 24.03 57.95 64.72 74.35 31.76 55.07
Uni-ControlNet (Zhao et al., 2024) all 32.58 78.08 29.37 65.85 44.35 77.40 37.98 66.75
UniControl (Qin et al., 2023) all 25.15 74.09 35.58 72.05 30.95 72.96 47.31 67.50
Cocktail (Hu et al., 2023) pose+hed 24.67 79.87 24.14 67.13 51.51 79.38 29.59 64.16
Ours all 12.01 81.36 43.85 74.11 11.51 81.66 60.22 69.28
Comparison of Image Quality and CLIP Score. To ascertain whether enhanced controllability correlates with a reduction in image quality, we present the Fréchet Inception Distance (FID) metrics across multiple conditional generation tasks, as detailed in Tab. 3. We can find that our model quantitatively reveals superior performance across all conditions compared to existing approaches. This significant achievement indicates that DynamicControl effectively handles intricate combinations of multiple spatial conditions, producing high-quality, coherent outcomes that align well with the spatial conditions. Further, to address concerns about its impact on text controllability, we employ CLIP-Score metrics to evaluate different methods across various datasets, ensuring that the generated images closely match the input text. As shown in Tab. 3, DynamicControl achieves superior CLIP-Score results on several datasets relative to existing methods.
Table 2: Results of combining different condition types. A quantitative comparison with models controlled by single visual conditions and various combinations of visual conditions. “Source” refers to the original reference image.
MultiGen-20M
Methods Conditions FID ($\downarrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$)
Source - - - 69.30
ControlNet++ canny 17.69 36.69 65.67
ControlNet++ hed 13.93 42.12 71.22
ControlNet++ depth 17.56 27.79 71.23
Ours pose 29.68 27.88 58.62
Ours pose+depth 19.85 31.69 61.02
Ours pose+depth+hed 13.55 41.52 66.35
Ours all 12.01 43.85 74.11
This indicates that our approach not only significantly improves conditional controllability, but also maintains the original model’s capability to generate images from text.
Table 3: FID ($\downarrow$) / CLIP-score ($\uparrow$) comparison under different conditional controls and datasets. All the results are conducted on $\times512$ image resolution for fair comparisons. We generate four groups of images and report the average result to reduce random errors.
Seg. Mask Canny Hed Openpose Depth
Gligen (Li et al., 2023c) 33.02 / 31.12 - 18.89 / 31.77 - 28.65 / 31.26 18.36 / 31.75
T2I-Adapter (Mou et al., 2024) 39.15 / 30.65 - 15.96 / 31.71 - 26.07 / 33.54 22.52 / 31.46
UniControlNet (Zhao et al., 2024) 39.70 / 30.59 - 17.14 / 31.84 17.08 / 31.94 27.66 / 34.58 20.27 / 31.66
UniControl (Qin et al., 2023) 46.34 / 30.92 - 19.94 / 31.97 15.99 / 32.02 24.58 / 35.01 18.66 / 32.45
ControlNet (Zhang et al., 2023) 33.28 / 31.53 21.33 / 13.31 14.73 / 32.15 15.41 / 32.33 - 17.76 / 32.45
Cocktail (Hu et al., 2023) 31.56 / 31.77 19.35 / 13.68 12.92 / 33.16 14.71 / 33.07 22.59 / 35.78 -
ControlNet++ (Li et al., 2025) 29.49 / 31.96 19.29 / 13.13 18.23 / 31.87 15.01 / 32.05 - 16.66 / 32.09
Figure 5: Comparison of multi-condition image generation by DynamicControl with different combinations against single-condition image generation by ControlNet++.
Condition Types Comparison. To examine the effect of the number of condition types, Fig.5 shows the visual results under various condition combinations, with qualitative evaluation metrics provided in Tab.2. Examples in Fig.5 demonstrate that as the number of conditions increases, the layout and texture within images become increasingly refined and accurate, demonstrating that the control effects of different visual conditions are not identical. Furthermore, as can be seen from Tab.2, an increase in visual conditions is beneficial for the overall quality and controllability of the images, which aligns with the observed visual effects. As evidenced by the comparison with ControlNet++ (Li et al., 2025), integrating multi-conditions improves control precision and enables finer-grained manipulation of the generated output.
Qualitative Comparison In Fig. 4, we visually compare different tasks in both single and multiple conditions. Our method consistently outperforms other models in terms of both visual quality and alignment with the specified conditions or prompts. This demonstrates the effectiveness of our approach in handling a diverse range of image generation tasks, while maintaining high fidelity to the input conditions. The comparative visual analysis highlights the robustness and adaptability of our method, demonstrating its efficacy across a broad spectrum of tasks.
4.2 Ablation Study
Loss Functions. To assess the effectiveness of different loss functions, we start with the base model which only contains the diffusion training loss. As shown in Sec. 4.2, adding the loss from the double-cycle controller significantly improves the alignment between the generated images and instruction conditions, as evidenced by the improved scores. Finally, the incorporation of combining the losses from LLM yields the best results across all metrics, underscoring the significance of condition evaluator for achieving high-fidelity and semantically accurate multiple condition generation.
Table 4: Ablation on loss functions. CLIP score and FID are reported on ADE20K and MultiGen-20M datasets. Base is a custom reduced version of DynamicControl.
ADE20K MultiGen-20M
Base 28.18 38.76 31.26 22.58
$\mathcal{L}_{\text{condition}}$ 30.11 33.59 33.05 19.62
$\mathcal{L}_{\text{image}}$ 31.23 29.66 34.22 18.22
$\mathcal{L}_{\text{LLM}}$ 33.04 27.58 35.26 16.85
$\mathcal{L}_{\text{eval}}$ 34.38 25.23 36.68 15.28
Figure 7: Results of multiple conditions correspond to different subjects.
Selection Schemes. As mentioned in Sec. 3.3, one of our key designs is to select dynamic number of conditions when performing the multi-control adapter, where we have tried different strategies of fixed numbers and adaptive number iteration. As shown in Sec. 4.2, in the scheme with a fixed number of iterations, the results vary with the iteration, but are still lower than those in the adaptive iteration method. This largely illustrates the effectiveness of the dynamic condition selection method in the multiple condition generation task.
Limitation Discussion. As shown in Fig. 7, we illustrate the results of multiple mismatched conditions, which demonstrates that our method has learned the combination of different conditions. But we still want to clarify that this non-aligned control generation is completely different from our setting. What we want to address is the complexity of multiple conditions and their potential conflicts for the same subject as we describe in the abstract and introduction. And we will dive into the exploration of this non-aligned control generation task in the future.
5 Conclusion
In this paper, we demonstrate from both quantitative and qualitative perspectives that existing works focusing on controllable generation still fail to fully harness the potential of multiple control conditions, leading to inconsistency between generated images and input conditions. To address this issue, we introduce DynamicControl , it explicitly optimizes the consistency between multiple input conditions and generated images using an efficient condition evaluator to rank the conditions, which integrates MLLM’s reasoning capabilities into the T2I generation task. Experimental results from various conditional controls reveal that DynamicControl substantially enhances controllability, without sacrificing image quality or image-text alignment. This provides fresh perspectives on controllable visual generation.
Appendix
A Reproducibility Statement
We have already elaborated on all the models or algorithms proposed, experimental configurations, and benchmarks used in the experiments in the main body or appendix of this paper. Furthermore, we declare that the entire code used in this work will be released after acceptance.
B The Use of Large Language Models
We use large language models solely for polishing our writing, and we have conducted a careful check, taking full responsibility for all content in this work.
C Implementation Details
Datasets. In our framework, we include up to 12 control conditions, the majority of which are sourced from the MultiGen-20M dataset, as proposed by UniControl (Qin et al., 2023). This dataset is a specialized subset derived from the larger LAION-Aesthetics (Schuhmann et al., 2022). More specifically, for the segmentation mask condition, our framework utilizes the ADE20K (Zhou et al., 2017; 2019) and COCOStuff (Caesar et al., 2018) datasets, following ControlNet (Zhang et al., 2023). For instances where the text caption data is missing in ADE20K, we supplement it with data from ControlNet++ (Li et al., 2025). In the testing dataset, due to the small size of the MultiGen-20M test set, and to ensure fairness in comparison, we combine the test and validation sets of MultiGen-20M to form a test set of up to 5500 text-image pairs. Additionally, to increase the confidence in our conclusions, we randomly sample 5000 text-image pairs from the Subject-200K(Tan et al., 2024) dataset as a second test set.
Evaluation Metrics. Our DynamicControl is trained using the training subsets of the respective datasets, and evaluations of all methods are conducted on the validation subsets. To ensure a fair comparison, the resolution for both training and inference in our framework is set at $\times512$ for all datasets and methods involved. Since the existing methods do not have so many control conditions, in order to facilitate fair comparison, we compare the quantitative evaluation on five conditions with more common control conditions: canny, hed, segmentation mask, openpose and depth. For each condition, controllability is assessed by quantifying the resemblance between the input conditions and the conditions extracted from the images generated by diffusion models. For the evaluation of segmentation, openpose and depth controls, we employ the mIoU, mAP and RMSE as metrics respectively, which is a common practice in related research fields. In the task of canny detection, we utilize the F1-Score as it effectively addresses the binary classification of pixels into categories of 0 (non-edge) and 1 (edge). This metric is particularly suitable given the pronounced long-tail distribution observed in edge data (Xie & Tu, 2015). For evaluating other methods, we use the open-source code provided by the respective developers to generate images. We ensure a fair comparison by conducting evaluations under identical conditions, using the same datasets and without altering their inference configurations.
Details of Multi-control Adapter. As illustrated in Fig. 2, these selected conditions are then consumed by Mixture-of-Experts (MoEs) (Shazeer et al., 2017), where all conditions are captured in parallel as features of various low-level visual conditions. Subsequently, the extracted features are channeled into $L$ blocks, each comprising criss-cross attention (Huang et al., 2019) and cross attention mechanisms. Within each block, the input features are initially processed through $n(n\leq N)$ criss-cross attentions operating in parallel, aligning them across various feature space dimensions. Following this alignment, the diversified features are concatenated, which are then forwarded to the cross attention module. This sequential processing ensures a comprehensive integration of feature dimensions, enhancing the depth and relevance of the attention mechanisms applied. Finally, these multi-control embeddings are utilized to direct the generative process, ensuring that the output is aligned with the nuanced requirements specified by the multiple control conditions.
Baselines. Our evaluation primarily focuses on several leading methods in the realm of controllable text-to-image diffusion models, including Gligen (Li et al., 2023c), T2I-Adapter (Mou et al., 2024), ControlNet v1.1 (Zhang et al., 2023), GLIGEN (Li et al., 2023c), Uni-ControlNet (Zhao et al., 2024), UniControl (Qin et al., 2023), Cocktail (Hu et al., 2023) and ControlNet++ (Li et al., 2025). These methods are pioneering in their field and provide public access to their codes and model weights, which accommodate various image conditions. Other approaches such as AnyControl (Sun et al., 2024), although their models are public, their code cannot be successfully run after many attempts.
Training Details. During the first stage of training, we adopt the pre-trained LLaVAv1.1-7B (Liu et al., 2024b) and QFormer (Li et al., 2023a) and employ DeepSpeed (Aminabadi et al., 2022) Zero-2 to perform LoRA (Hu et al., 2021) fine-tuning. The Stable Diffusion-1.5 (Rombach et al., 2022) is diffusion model pre-trained weights. The learning rate and weight decay parameters are set to 2e-4 and 0, respectively. In the second stage, the values of learning rate, weight decay, and warm-up ratio are set to 1e-5, 0, and 0.001, respectively. We train the model for totally 50K iterations. Furthermore, we take the AdamW (Loshchilov, 2017) as the optimizer based on PyTorch Lightning (Paszke et al., 2019) for both stages. Our full-version DynamicControl is trained on 8 H20 GPUs with the batch size of 4 and the whole training process can be completed within 3 days. It takes an average of 34s to generate one image on one H20 GPU. The values of $\lambda$ in Eq.10 are experimentally tested through a series of experiments, which are set to 2 and 1.5 respectively in our practice. These values vary from 0.1 to 3 at every 0.1 interval. In all the experiments, we adopt DDIM (Song et al., 2020a) sampler with 50 timesteps for all the compared methods.
D More Experiments
D.1 Comparison of Controllability
[…]
@@ D.1 Comparison of Controllability
Dataset
T2I
Model
MultiGen-20M MultiGen-20M MultiGen-20M MultiGen-20M ADE20K COCO-Stuff
ControlNet (Zhang et al., 2023) SDXL - - - 40.01 - -
T2I-Adapter (Mou et al., 2024) SDXL 28.03 - 63.89 39.76 - -
T2I-Adapter (Mou et al., 2024) SD1.5 23.66 - 60.17 48.40 12.60 -
Gligen (Li et al., 2023c) SD1.4 26.92 0.5641 69.88 38.82 23.77 -
Uni-ControlNet (Zhao et al., 2024) SD1.5 27.31 0.6912 72.71 40.66 19.39 -
UniControl (Qin et al., 2023) SD1.5 30.83 0.7967 75.87 39.17 25.45 -
ControlNet (Zhang et al., 2023) SD1.5 34.66 0.7622 - - 32.56 27.47
Cocktail (Hu et al., 2023) SD1.5 35.22 0.8152 78.82 35.90 36.55 29.68
ControlNet++ (Li et al., 2025) SD1.5 37.04 0.8097 - 28.32 43.64 34.56
@@
Table A1: Controllability comparison under different conditional controls and datasets. We generate four groups of images and report the average result to reduce random errors.
D.2 Comparison on Computational Complexities